OpenStudy (anonymous):

What is the y coordinate of the y−intercept of the line that passes through the points (−3, −1) and (1, 7)? Numerical Answers Expected! ANSWER ONLY .

OpenStudy (anonymous):

y1-y2/x1-x2 -1-7/-3-1

OpenStudy (anonymous):

-8/-4 =2

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

y=mx+b 7=m(1)+b 7=2+b -2 -2+b 5=b

OpenStudy (anonymous):

y=intercept=5